OpenStudy (anonymous):

evaluate using the substitution method: 8/(x-9)^3 dy a. 2/(x-9)^4+C b. -4/(x-9)^2+C c. -2/(x-9)^2+C d. 4/(x-9)^2+C

OpenStudy (anonymous):

Pretty straight forward, using the formula: u = x - 9 --> du = .. Plug it in:

OpenStudy (anonymous):

I got A at first but I didn't know if it was correct or not...

OpenStudy (anonymous):

Sorry, try again!

OpenStudy (anonymous):

Can you show your steps here?

OpenStudy (anonymous):

u=x-9 du=dx \[\int\limits 8/(x-9)^3dx=8\int\limits1/u^3 du=-4/u^2+c=-4/(x-9)^2+c\]